An injection-molded part is likely to be obtained from any one of 8 cavities on a mold. Round your answers to three decimal places (e.g. 98.765). (b) What is the probability that a part is from cavity 1, 2 or 3? 0.33 (c) 50.77
Answer #1

b)

since there are total 8 cavity and 1,2 or 3 is 3 out of them

probability =3/8 =0.375

c)

P(neither from 3 nor 4 )=P(from other 6) =6/8 =0.75